OpenStudy (anonymous):

Identify the kind of transformation from g(x)=x4 to h(x)=(6x)4

OpenStudy (anonymous):

x4=x^4 , (6x)4 = (6x)^4 from (goh(x)=((6x)^4)^4 =(6x)^(4*4) =(6x)^16